How many cubed pieces of fudge that are 3 inches on an edge can be packed into a tin that is 9 inches deep by 12 inches wide by 8 inches high with the lid still being able to close
Answer by london maths tutor(243)

Volume of the fudge = 3*3*3 = 27 inches^3
Volume of the tin = 9*12*8 = 864 inches^3
Number of cubes of fudge = 864/27 = 32
answer: 32 pieces of fudge.